# A Majorization-Minimization Gauss-Newton Method for 1-Bit Matrix Completion

### v2

@@ 4.2.4 Varying the matrix rank
In the last simulation under the non-spiky case, we examined how the performance of each method scales with the rank of the underlying matrix. We fixed the size of ${\bm{\mathbf{{\Theta}}}}^{*}$ at $m=n=1000$ and generated matrices of rank $r^{*}\in\{3,5,8,10\}$. The average spikiness ratios of the generated underlying matrices for the corresponding values of $r^{*}$ were $\{4.54,4.81,5.03,4.97\}$. We considered the probit model with a noise level $\sigma=0.18$ and set the observation fraction $\rho=0.8$. For all methods except for MaxNorm, we estimated the rank $r$ from $\{1,2,\cdots,12\}$ using the validation procedure in Section 4.1.
Figure 4: Probit model ($\sigma=0.18$): Median over 20 replicates of relative error, Hellinger distance, and runtime versus true rank $r^{*}$ for a non-spiky underlying matrix of size $m\times n=1000\times 1000$ and observation fraction $\rho=0.8$.
Figure 5: Probit model ($\sigma=0.18$): Median over 20 replicates of relative error (using the training set), log-likelihood (on the testing set), and runtime versus estimated rank $\hat{r}$ for a non-spiky underlying matrix of size $m\times n=1000\times 1000$, true rank $r^{*}=5$, and observation fraction $\rho=0.8$.
Figure 4 displays the performance of different methods under varying rank $r^{*}$. As expected, as $r^{*}$ increased, the estimation error of each method increased since the number of parameters to be estimated increased. All methods behaved comparably in estimation accuracy, with the Hellinger distance of TraceNorm slightly worse than the other methods across different ranks. MMGN enjoyed the fastest runtime among all the compared methods. Recall that we provided the true rank $r^{*}$ for MaxNorm, thus its computational speed would be even longer if the rank were selected by validation.
To evaluate the rank estimation procedure and illustrate how different methods perform given different estimated ranks, we reran the above simulation with the true rank $r^{*}=5$ and reported the relative error using the training set, the likelihood on the testing set, and the runtime of each method when given an estimated rank $\hat{r}$ from the set $\{1,2,\cdots,12\}$. Note that MaxNorm was not included since we ran it without the rank estimation procedure. As shown in Figure 5, all methods achieved the highest log-likelihood as well as the lowest estimation error at the true rank $r^{*}=5$. In other words, they selected the correct rank by the procedure. As shown in the right panel of Figure 5, MMGN, GD, LBFGS, and Manopt required longer runtimes when given a rank higher than the true rank. This explains why the runtimes of the four methods in Figure 4 decreased as the true rank $r^{*}$ increased. With a fixed set of $\hat{r}\in\{1,2,\cdots,12\}$, as $r^{*}$ increased from $3$ to $10$, the number of values of $\hat{r}$ that were higher than $r^{*}$ decreased, which led to shorter overall runtime.
@@ 4.3 Spiky Matrices
Figure 6: Probit model ($\sigma=2$): Median over 20 replicates of relative error, Hellinger distance, and runtime versus observation fraction $\rho$ at low (top), intermediate (middle), and high (bottom) spikiness levels for an underlying matrix of size $m\times n=1000\times 1000$ and rank $r^{*}=1$. Average spikiness ratios are $17.57$ (top), $33.16$ (middle), and $47.19$ (bottom).
@@
An interesting phenomenon is that for all methods except for TraceNorm, their errors in estimating ${\bm{\mathbf{{\Theta}}}}^{*}$ became relatively insensitive to $\rho$ as the spikiness level increased. The relative errors were dominated by errors in the large-magnitude entries, and errors in these large-magnitude entries were hardly affected by $\rho$. These methods produced accurate estimates for the majority of low-magnitude entries, even when $\rho$ was small. Therefore, increasing $\rho$ did not lead to significant improvement in their estimation errors. In contrast, TraceNorm required a large value of $\rho$ to produce good estimates for low-magnitude entries. As a result, we saw a more noticeable decrease in its estimation error as $\rho$ grew. We present a detailed example to elaborate on this phenomenon in the supplementary materials.
The right panel of each row in Figure 6 shows the computational advantage of MMGN when dealing with spiky matrices at different spikiness levels. It ran dozens of times faster than TraceNorm and MaxNorm. Compared to the three generic schemes for the optimization problem (2), MMGN ran typically two to ten times faster. Only LBFGS exhibited runtimes comparable with MMGN but only when the observation fraction $\rho$ was close to one.
@@ 4.4 Real Data Example
We applied MMGN to the MovieLens (1M) data set [^2] The data set is available at http://www.grouplens.org/node/73., which is larger but otherwise similar to the MovieLens (100K) real data example in Davenport et al. (2014). It contains $1{,}000{,}209$ movie ratings from $6{,}040$ users on $3{,}952$ movies, with each rating on a scale from $1$ to $5$. The resulting observation fraction is $0.04$. Following Davenport et al. (2014), we converted all ratings to binary observations by comparing each rating to the average rating over all movies. Ratings above the average were encoded as $+1$ and $-1$ otherwise. As in Davenport et al. (2014), we considered the logistic model with the noise level $\sigma=1$ and used $95\%$ of the ratings as the training set to estimate ${\bm{\mathbf{{\Theta}}}}^{*}$. The performance was evaluated on the remaining $5\%$ of ratings by checking whether or not the estimate of ${\bm{\mathbf{{\Theta}}}}^{*}$ correctly predicts the sign of the ratings (above or below the average rating). As in our simulations, we compared the performance of MMGN to TraceNorm, MaxNorm, GD, LBFGS, and Manopt. We followed Davenport et al. (2014) and treated $\alpha\beta$ as a single parameter in TraceNorm. We assigned $\alpha\beta$ ten values equally spaced on a logarithmic scale between $10^{-0.5}$ to $10$. We ran MMGN and the three generic schemes with a candidate rank $r$ ranging from $1$ to $10$. For MaxNorm, $\beta$ is set to be $\sqrt{r}$, where again $r$ ranged from $1$ to $10$, and $\alpha$ is chosen such that $\alpha\beta$ takes the same value as $\alpha\beta$ for TraceNorm. As in Davenport et al. (2014), we selected parameter values for each method that led to the best prediction performance on the separate validation set.
@@ 4.5 Additional Comparisons
As mentioned above, 1-bit matrix completion can be viewed as a logistic PCA problem but with only partially observed entries. In particular, one could solve the 1-bit objective by a logistic-PCA scheme that is able to handle missing data. Indeed, De Leeuw (2006) proposed such a scheme, denoted logisticPCA. Thus, we compared MMGM to logisticPCA. Interestingly, this scheme is also based on MM. The difference between the two algorithms lies in how they solve the rank-constrained least squares problem. The per-iteration computation in logisticPCA is dominated by a rank-$r$ truncated SVD. The implementation we used for logisticPCA is based on ARPACK which requires $r$ rounds of the Lanczos method for a total cost of $O(mnr)$ flops; each round of the Lanczos method costs $O(mn)$ flops, due to matrix-vector multiplies involving an $m$-by-$n$ matrix (Lehoucq et al., 1998). In contrast, MMGN adopts a single Gauss-Newton step to approximate the rank-constrained least squares problem by a sparse least square problem whose computational cost is $\mathcal{O}(|\Omega|r)$. A detailed breakdown of this cost is provided in the supplementary materials. In simulations, which can be found in the supplementary materials, logisticPCA typically achieved accuracies comparable with MMGN but ran two to ten times slower than MMGN, especially in the more challenging cases where the observation fraction is small.
@@ 5 Summary and Discussion
In this paper, we proposed MMGN, a novel fast and accurate algorithm for solving the optimization problem. Our algorithm employs the MM principle to convert the original challenging problem (2) into a sequence of standard low-rank matrix completion problems. For each MM update, we apply a factorization strategy to incorporate the low-rank structure, resulting in a nonlinear least squares problem. We then apply a modified Gauss-Newton scheme to compute an inexact MM update by solving a linear least squares problem. Hence, in comparison to previous works, our method is relatively simple and can easily scale to large matrices. In comparison to two established 1-bit matrix completion methods, TraceNorm (Davenport et al., 2014) and MaxNorm (Cai and Zhou, 2013), and three generic schemes for solving the original optimization problem (2), MMGN typically achieved comparable and sometimes lower errors while often being significantly faster.
@@ Acknowledgments
B.N. is the incumbent of the William Petschek Professorial Chair of Mathematics. The research of B.N. was funded by the National Institutes of Health (NIH) grant R01GM135928 and by the Israel Science Foundation (ISF) grant 2362/22. The research of E.C. was supported by the NIH grant R01GM135928. We thank Wenxin Zhou for sharing us the code of MaxNorm. We thank Mark Davenport for helpful discussions.

### v3

@@ 4.2.4 Varying the matrix rank
In the last simulation under the non-spiky case, we examined how the performance of each method scaled with the rank of the underlying matrix. We fixed the size of ${\bm{\mathbf{{\Theta}}}}^{*}$ at $m=n=1000$ and generated matrices of rank $r^{*}\in\{3,5,8,10\}$. The average spikiness ratios of the generated underlying matrices for the corresponding values of $r^{*}$ were $\{4.54,4.81,5.03,4.97\}$. We considered the probit model with a noise level $\sigma=0.18$ and set the observation fraction $\rho=0.8$. For all methods except for MaxNorm, we estimated the rank $r$ from $\{1,2,\cdots,12\}$.
Figure 4: Probit model ($\sigma=0.18$): Relative error, Hellinger distance, and runtime versus true rank $r^{*}$ for a non-spiky underlying matrix of size $m\times n=1000\times 1000$ and observation fraction $\rho=0.8$.
Figure 5: Probit model ($\sigma=0.18$): Relative error (using the training set), log-likelihood (on the testing set), and runtime versus estimated rank $\hat{r}$ for a non-spiky underlying matrix of size $m\times n=1000\times 1000$, true rank $r^{*}=5$, and observation fraction $\rho=0.8$. LBFGS produced log-likelihood of negative infinity at $\hat{r}\in\{11,12\}$, so the corresponding box plots were missing.
Figure 4 displays the performance of different methods under varying rank $r^{*}$. As expected, as $r^{*}$ increased, the estimation error of each method increased since the number of parameters to be estimated increased. All methods behaved comparably in estimation accuracy, with the Hellinger distance of TraceNorm slightly worse than the other methods across different ranks. MMGN enjoyed the fastest runtime among all the compared methods. Recall that we provided the true rank $r^{*}$ for MaxNorm, thus its runtime would be even longer if the rank were selected by validation.
To evaluate the rank estimation procedure and illustrate how different methods perform given different estimated ranks, we reran the above simulation with the true rank $r^{*}=5$ and reported the relative error using the training set, the log-likelihood on the testing set, and the runtime of each method when given an estimated rank $\hat{r}$ from the set $\{1,2,\cdots,12\}$. Note that MaxNorm was not included since we ran it without the rank estimation procedure. As shown in Figure 5, all methods achieved the highest log-likelihood as well as the lowest estimation error at the true rank $r^{*}=5$. In other words, they selected the correct rank by the procedure. As shown in the right panel of Figure 5, MMGN, GD, LBFGS, and Manopt required longer runtimes when given a rank higher than the true rank. This explains why the runtimes of the four methods in Figure 4 decreased as the true rank $r^{*}$ increased. With a fixed set of $\hat{r}\in\{1,2,\cdots,12\}$, as $r^{*}$ increased from $3$ to $10$, the number of values of $\hat{r}$ that were higher than $r^{*}$ decreased, which led to shorter overall runtime.
@@ 4.3 Spiky Matrices
Figure 6: Probit model ($\sigma=2$): Relative error, Hellinger distance, and runtime versus observation fraction $\rho$ at low (top), intermediate (middle), and high (bottom) spikiness levels for an underlying matrix of size $m\times n=1000\times 1000$ and rank $r^{*}=1$. Average spikiness ratios are $17.57$ (top), $33.16$ (middle), and $47.19$ (bottom).
@@
An interesting phenomenon is that for all methods except TraceNorm, their errors in estimating ${\bm{\mathbf{{\Theta}}}}^{*}$ became relatively insensitive to $\rho$ as the spikiness level increased. The relative errors were dominated by errors in the large-magnitude entries, and errors in these entries were mildly affected by $\rho$. These methods produced accurate estimates for the majority of low-magnitude entries, even when $\rho$ was small. Therefore, increasing $\rho$ did not lead to significant improvement in their estimation errors. In contrast, TraceNorm required a large value of $\rho$ to produce good estimates for low-magnitude entries. As a result, we saw a more noticeable decrease in its estimation error as $\rho$ grew. We present a detailed example to elaborate on this phenomenon in the supplementary materials.
The right panel of each row in Figure 6 shows the computational advantage of MMGN when dealing with spiky matrices at different spikiness levels. It ran dozens of times faster than TraceNorm and MaxNorm. Compared to the three generic schemes for the optimization problem (2), MMGN ran typically two to ten times faster. LBFGS exhibited runtimes comparable with MMGN but only when the observation fraction $\rho$ was close to one.
@@ 4.4 Real Data Example
We applied MMGN to the MovieLens (1M) data set [^2] The data set is available at http://www.grouplens.org/node/73., which is larger but otherwise similar to the MovieLens (100K) real data example in Davenport et al. (2014). It contains $1{,}000{,}209$ movie ratings from $6{,}040$ users on $3{,}952$ movies, with each rating on a scale from $1$ to $5$. The resulting observation fraction is $0.04$. Following Davenport et al. (2014), we converted all ratings to binary observations by comparing each rating to the average rating over all movies. Ratings above the average were encoded as $+1$ and $-1$ otherwise. As in Davenport et al. (2014), we considered the logistic model with the noise level $\sigma=1$ and used $95\%$ of the ratings as the training set to estimate ${\bm{\mathbf{{\Theta}}}}^{*}$. The performance was evaluated on the remaining $5\%$ of ratings by checking whether or not the estimate of ${\bm{\mathbf{{\Theta}}}}^{*}$ correctly predicts the sign of the ratings (above or below the average rating). We compared MMGN to TraceNorm, MaxNorm, GD, LBFGS, and Manopt. We followed Davenport et al. (2014) and treated $\alpha\beta$ as a single parameter in TraceNorm. We assigned $\alpha\beta$ ten values equally spaced on a logarithmic scale between $10^{-0.5}$ to $10$. We ran MMGN and the three generic schemes with a candidate rank $r$ ranging from $1$ to $10$. For MaxNorm, $\beta$ was set to $\sqrt{r}$, where again $r$ ranged from $1$ to $10$, and $\alpha$ was chosen such that $\alpha\beta$ took the same value as $\alpha\beta$ for TraceNorm. As in Davenport et al. (2014), we selected parameter values for each method that led to the best prediction performance on the separate validation set.
@@ 4.5 Additional Comparisons
As mentioned above, 1-bit matrix completion can be viewed as a logistic PCA problem but with only partially observed entries. In particular, one could solve the 1-bit objective by a logistic-PCA scheme that is able to handle missing data. De Leeuw (2006) proposed such a scheme, denoted logisticPCA. Thus, we compared MMGM to logisticPCA. Interestingly, this scheme is also based on MM. The difference between the two algorithms lies in how they solve the rank-constrained least squares problem. The per-iteration computation in logisticPCA is dominated by a rank-$r$ truncated SVD. The implementation we used for logisticPCA is based on ARPACK which requires $r$ rounds of the Lanczos method for a total cost of $O(mnr)$ flops; each round of the Lanczos method costs $O(mn)$ flops, due to matrix-vector multiplies involving an $m$-by-$n$ matrix (Lehoucq et al., 1998). In contrast, MMGN adopts a single Gauss-Newton step to approximate the rank-constrained least squares problem by a sparse least square problem whose computational cost is $\mathcal{O}(|\Omega|r)$. A detailed breakdown of this cost is provided in the supplementary materials. In simulations, which can be found in the supplementary materials, logisticPCA typically achieved accuracies comparable with MMGN but ran two to ten times slower than MMGN, especially in the more challenging cases where the observation fraction is small.
@@ 5 Summary and Discussion
In this paper, we proposed MMGN, a novel fast and accurate algorithm for solving the 1-bit optimization problem. Our algorithm employs the MM principle to convert the original challenging problem (2) into a sequence of standard low-rank matrix completion problems. For each MM update, we apply a factorization strategy to incorporate the low-rank structure, resulting in a nonlinear least squares problem. We then apply a modified Gauss-Newton scheme to compute an inexact MM update by solving a linear least squares problem. Hence, in comparison to previous works, our method is relatively simple and can easily scale to large matrices. In comparison to two established 1-bit matrix completion methods, TraceNorm (Davenport et al., 2014) and MaxNorm (Cai and Zhou, 2013), and three generic schemes for solving the original optimization problem (2), MMGN typically achieved comparable and sometimes lower errors while often being significantly faster.
@@ Acknowledgments
B.N. is incumbent of the William Petschek Professorial Chair of Mathematics. The research of B.N. was funded by the National Institutes of Health (NIH) grant R01GM135928 and by the Israel Science Foundation (ISF) grant 2362/22. The research of E.C. was supported by the NIH grant R01GM135928. We thank Wenxin Zhou for sharing us the code of MaxNorm. We thank Mark Davenport for helpful discussions.